Q. If x=a(@+sin@) and y=a(1-cos@), find dy/dx.

Grade:11

2 Answers

honey trivedi
13 years ago

i think the correct ans is 1/a

Chetan Mandayam Nayakar
13 years ago

dy/da = 1-cos@ + @sin@

dx/da= @+sin@ +a(1+cos@) = acos@ +sin@ +2a

dy/dx=(dy/da)/(dx/da) = (asin@ -cos@ +1)/(acos@ + sin@ +2a)